i need help...graph the equation using x and y intercepts...
x/2-y/5=1... how do i graph this?

Answer by mananth(16949)

x/2-y/5=1
LCD = 10
mutiply equation by LCD
5x-2y=10
2y=5x-10
y= (5/2)x-5
Generate values of x & y to draw the graph
x=2, y=0 (2,0)
x=-2, y=-10, (-2,-10)
x=4,y=5 , (4,5)